# Learning a Gaussian Mixture for Sparsity Regularization in Inverse Problems

### v1

@@ 6.3. Denoising
In Table 1 we show the mean over $2000$ signals of the test set of the relative MSE between the original signals and the different reconstructions. We observe that our unsupervised algorithm (B) provides the best reconstructions for Datasets 1 and 3, and for Dataset 2 it is only outperformed by group dictionary learning (D). However, dictionary learning is computationally much more expensive than our unsupervised method, and the gain in terms of reconstruction is minimal (as can be visually seen in Figure 2). From Table 1 it seems clear that the best baseline algorithms to which we can compare our approaches are dictionary learning (C), among the methods that do not use clustering, and group dictionary learning (D), among the methods using the groups obtained by the clustering of the training set. For this reason, in Figure 2 we show a qualitative comparison between the reconstructions provided by our methods, dictionary learning, and group dictionary learning. For completeness, we provide the remaining qualitative comparisons in Figure 4 of Appendix C. We notice that for Datasets 2 and 3 the algorithms learn all the discontinuities and try to remove those not present in the signal.
@@
Supervised (A) $1.97\%$ $7.03\hskip 2.84544pt10^{-3}\%$ $2.71\hskip 2.84544pt10^{-2}\%$
Unsupervised (B) $\mathbf{0.98}\%$ $1.80\hskip 2.84544pt10^{-3}\%$ $\mathbf{6.32\hskip 2.84544pt10^{-3}}\%$
Dictionary learning (C) $4.18\%$ $3.43\hskip 2.84544pt10^{-3}\%$ $7.13\hskip 2.84544pt10^{-3}\%$
Group dictionary learning (D) $0.99\%$ $\mathbf{1.70\hskip 2.84544pt10^{-3}}\%$ $2.42\hskip 2.84544pt10^{-2}\%$
IHT with SVD basis (E) $9.93\%$ $1.25\hskip 2.84544pt10^{-2}\%$ $3.97\hskip 2.84544pt10^{-1}\%$
IHT with SVD bases of groups (F) $0.99\%$ $2.04\hskip 2.84544pt10^{-3}\%$ $3.97\hskip 2.84544pt10^{-1}\%$
IHT with known basis (G) $2.78\%$ $1.22\hskip 2.84544pt10^{-2}\%$ $2.46\hskip 2.84544pt10^{-2}\%$
LASSO with SVD basis (H) $9.24\%$ $1.55\hskip 2.84544pt10^{-2}\%$ $3.07\hskip 2.84544pt10^{-1}\%$
Group LASSO with SVD bases (I) $3.01\%$ $3.71\hskip 2.84544pt10^{-3}\%$ $8.31\hskip 2.84544pt10^{-1}\%$
LASSO with known basis (J) $4.69\%$ $1.00\hskip 2.84544pt10^{-2}\%$ $2.15\hskip 2.84544pt10^{-2}\%$
Figure 2. Qualitative comparison for the denoising problem. In each column we show a signal of the test set from Datasets 1, 2 and 3, respectively. In each row we report the original data in orange and the noisy data, the reconstructions obtained with the methods supervised (A), unsupervised (B), dictionary learning (C), and group dictionary learning (D), respectively, in blue.
@@
Exact $0.97\%$ $1.66\hskip 2.84544pt10^{-3}\%$ $3.37\hskip 2.84544pt10^{-3}\%$
Learned $0.98\%$ $1.80\hskip 2.84544pt10^{-3}\%$ $6.32\hskip 2.84544pt10^{-3}\%$
Random $8.25\pm 0.04\%$ $(3.46\pm 0.25)\hskip 2.84544pt10^{-3}\%$ $(5.70\pm 0.55)\hskip 2.84544pt10^{-3}\%$
@@ 6.4. Deblurring
Unsupervised (B) $\mathbf{3.68}\%$ $\mathbf{2.65\hskip 2.84544pt10^{-3}}\%$ $\mathbf{1.01\hskip 2.84544pt10^{-2}}\%$
Dictionary learning (C) $14.32\%$ $6.61\hskip 2.84544pt10^{-3}\%$ $1.28\hskip 2.84544pt10^{-2}\%$
Group dictionary learning (D) $13.51\%$ $4.62\hskip 2.84544pt10^{-3}\%$ $3.41\hskip 2.84544pt10^{-2}\%$
IHT with SVD bases of groups (F) $3.80\%$ $5.54\hskip 2.84544pt10^{-3}\%$ $9.48\hskip 2.84544pt10^{-1}\%$
Group LASSO with SVD bases (I) $11.48\%$ $1.34\hskip 2.84544pt10^{-2}\%$ $9.11\hskip 2.84544pt10^{-1}\%$
@@ 7. Conclusions
To validate our approach, we conducted numerical implementations (both supervised and unsupervised) and compared them against established baseline algorithms for sparse optimization, such as Lasso, IHT, and their group variants. Additionally, we incorporated a dictionary learning strategy for fair comparisons. The experiments focused on three 1D datasets featuring sparse or compressible signals, addressing denoising and deblurring tasks.
@@ Acknowledgements
Then, using [27, Theorem $2$, Section $13$], the conditional distribution $(X_{i}|Y_{i}=y)$ is Gaussian and its expectation is given by
@@ B.2. Group LASSO
The minimization of (18) can be performed via the iterative algorithm proposed in [30, Proposition $1$], showing strong connections with the proximal gradient descent method, which reads
@@ Appendix C Additional numerical results
In Figure 4, we show a qualitative comparison for the denoising problem described in Section 6.3 between the reconstructions obtained with our unsupervised method, IHT with SVD basis, IHT with SVD bases of groups, IHT with known basis, LASSO with SVD basis, Group LASSO with SVD bases and LASSO with known basis.
Figure 4. Further qualitative comparisons for the denoising problem. In each column we show a signal of the test set from Datasets 1, 2 and 3, respectively. In each row we report the original data in orange and the noisy data, and the reconstructions obtained with the unsupervised (B), IHT with SVD basis (E), IHT with SVD bases of groups (F), IHT with known basis (G), LASSO with SVD basis (H), Group LASSO with SVD bases (I), and LASSO with known basis (J) approach, respectively, in blue.

### v2

@@ 6.3. Denoising
In Table 1 we show the mean over $2000$ signals of the test set of the relative MSE between the original signals and the different reconstructions. We observe that our unsupervised technique (B) provides the best reconstructions for Datasets 1 and 3, and for Dataset 2 it is only outperformed by group dictionary learning (D). However, dictionary learning is computationally more expensive than our unsupervised method. Indeed, although both techniques involve a preliminary clustering step, group dictionary learning also requires computing $L$ sparsifying dictionaries, i.e., solving $L$ optimization problems as (20). In our experiments, for each dictionary, the online method presented in Section 5.4 performed nearly $300$ iterations, with the cost of a single update depending on the size of both the dictionary and the sample, as discussed in Section 5.4. On the other hand, the unsupervised technique only requires estimating $L$ means and covariances prior to the application of (9), and the difference in performance is minimal (as it can be seen in Figure 2). From Table 1 it seems clear that the best baseline algorithms to which we can compare our approaches are dictionary learning (C), among the methods that do not use clustering, and group dictionary learning (D), among the methods using the groups obtained by the clustering of the training set. For this reason, in Figure 2 we show a qualitative comparison between the reconstructions provided by our methods, dictionary learning, and group dictionary learning. For completeness, we provide the remaining qualitative comparisons in Figure 4 of Appendix C. We notice that for Datasets 2 and 3 the algorithms learn all the discontinuities and try to remove those not present in the signal.
@@
[…]
Figure 2. Qualitative comparison for the denoising problem. In each column we show a signal of the test set from Datasets 1, 2 and 3, respectively. In each row we report the original data in orange and the noisy data, the reconstructions obtained with the methods supervised (initialized with the unsupervised one) (A+B), unsupervised (B), dictionary learning (C), and group dictionary learning (D), respectively, in blue.
@@
[…]
Random $8.25$\pm$0.04$\%$ $(3.46\pm 0.25)\times 10^{-3}\%$ $(5.70\pm 0.55)\times 10^{-3}\%$
Does regularization really matter?
We investigated the impact of the regularization term $\mathcal{J}$ added to the training loss in the supervised approach, as discussed in Section 4.1. Table 3 shows that, when applied to Dataset 1, the use of the Frobenius or the nuclear norms does not improve the results in terms of MSE. In both cases, the value of the regularization parameter is heuristically chosen. Notice that also the matrices $\Sigma_{i}$ trained without any penalty term naturally show a good degree of sparsity (quantified by the ratio between their average rank and the size of the signal).
Table 3. Denoising problem with $10\%$ noise for Dataset 1 with the supervised approach (B) without penalty, with a nuclear-norm penalty, and with a Frobenius-norm penalty. We compare the values of the relative MSE and the ratio between the average rank of the learned $\Sigma_{i}$ and the signal size $n$.
No regularization Nuclear norm Frobenius norm
Relative MSE $1.97\%$ $2.18\%$ $2.11\%$
Average rank of $\Sigma_{i}$ 0.41 $n$ 0.22 $n$ 0.92 $n$
@@ 6.4. Deblurring
[…]
@@ 7. Conclusions
To validate our approach, we conducted numerical implementations (both supervised and unsupervised) and compared them against established baseline algorithms for sparse optimization, such as Lasso, IHT, and their group variants. Additionally, we incorporated a dictionary learning strategy for fair comparisons. The experiments focused on three datasets of 1D signals (discretized as vectors in $\mathbb{R}^{n}$) featuring sparse or compressible signals, addressing denoising and deblurring tasks.
@@ Acknowledgements
Then, using [31, Theorem $2$, Section $13$], the conditional distribution $(X_{i}|Y_{i}=y)$ is Gaussian and its expectation is given by
@@ B.2. Group LASSO
The minimization of (18) can be performed via the iterative algorithm proposed in [35, Proposition $1$], showing strong connections with the proximal gradient descent method, which reads
@@ Appendix C Additional numerical results
In Figure 4, we show a qualitative comparison for the denoising problem described in Section 6.3 between the reconstructions obtained with our supervised method (with random initialization), our unsupervised method, IHT with SVD basis, IHT with SVD bases of groups, IHT with known basis, LASSO with SVD basis, Group LASSO with SVD bases and LASSO with known basis.
Figure 4. Further qualitative comparisons for the denoising problem. In each column we show a signal of the test set from Datasets 1, 2 and 3, respectively. In each row we report the original data in orange and the noisy data, and the reconstructions obtained with the randomly-initialized supervised (A), unsupervised (B), IHT with SVD basis (E), IHT with SVD bases of groups (F), IHT with known basis (G), LASSO with SVD basis (H), Group LASSO with SVD bases (I), and LASSO with known basis (J) approach, respectively, in blue.